# AdaCAD: Adaptively Decoding to Balance Conflicts between Contextual and Parametric Knowledge

### v1

@@ Abstract
Knowledge conflict arises from discrepancies between information in the context of a large language model (LLM) and the knowledge stored in its parameters. This can hurt performance when using standard decoding techniques, which tend to ignore the context. Existing test-time contrastive methods seek to address this by comparing the LLM’s output distribution with and without the context and adjust the model according to the contrast between them. However, we find that these methods frequently misjudge the degree of conflict and struggle to handle instances that vary in their amount of conflict, with static methods over-adjusting when conflict is absent. We propose a fine-grained, instance-level approach called AdaCAD, which dynamically infers the weight of adjustment based on the degree of conflict, as measured by the Jensen-Shannon divergence between distributions representing contextual and parametric knowledge. Our experiments across four models on six diverse question-answering (QA) datasets and three summarization tasks demonstrate that our training-free adaptive method consistently outperforms other decoding methods on QA, with average accuracy gains of $14.21\%$ (absolute) over a static contrastive baseline, and improves the factuality of summaries by 5.59 (AlignScore). Furthermore, our analysis shows that while decoding with contrastive baselines hurts performance when conflict is absent, AdaCAD mitigates these losses, making it more applicable to real-world datasets in which some examples have conflict and others do not. [^1] Our code is publicly available at: https://github.com/HanNight/AdaCAD.
@@
Figure 1: In cases of high knowledge conflict, greedy decoding fails to attend to the context, resulting in incorrect answers. Contrastive decoding allows the context to be incorporated, but must be done with care: in low-conflict cases, excessive contrast (e.g., by CAD with $\alpha=1$) can over-correct, resulting in incorrect outputs. AdaCAD dynamically adjusts the degree of contrast, allowing it to handle both high and low-conflict cases.
@@ 1 Introduction
We demonstrate AdaCAD’s effectiveness on a diverse range of tasks, covering question-answering (QA) and summarization, with six QA datasets (Natural Question (NQ; Kwiatkowski et al., 2019), NQ-Swap (Longpre et al., 2021), TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2023), HotpotQA (Yang et al., 2018), TabMWP (Lu et al., 2023)) and three summarization datasets (CNN-DM (See et al., 2017), XSum (Narayan et al., 2018), and TofuEval (Tang et al., 2024)). We test a range of base LLMs, examining Llama-2 (Touvron et al., 2023), Llama-3 (AI@Meta, 2024), and Mistral (Jiang et al., 2023). We consider different sizes of these models and also test both the base and instruction-tuned variants. Our results and analyses show that decoding with a uniform level of contrast benefits high-conflict scenarios but generally hurts performance, while the adaptive contrast of AdaCAD results in improvements across the board. Overall, AdaCAD generally achieves superior performance compared to the baselines, with an absolute gain of $14.21\%$ over CAD (a static baseline), $4.82\%$ over COIECD (Yuan et al., 2024, a method that classifies instances as conflicting or not), and $2.41\%$ over greedy decoding when averaged across models and QA datasets. On summarization, AdaCAD improves summary quality and factuality, with an average AlignScore (Zha et al., 2023) gain of 4.16 over greedy decoding, 2.19 over CAD, and 10.44 over COIECD.
@@
Furthermore, in our analyses, we explore why AdaCAD improves over the baselines. We first validate the hypothesis that AdaCAD is able to balance contextual and parametric knowledge by assigning lower weights to lower-conflict instances, testing each method on datasets designed to have high and low conflict, finding that AdaCAD’s inferred weight is much lower when there is no conflict. We also compare the amount by which CAD and AdaCAD adjust the base model’s distribution on examples with and without conflict, finding that while AdaCAD changes the distribution less when there is no conflict (i.e. when the base model’s distribution is already sufficient), CAD adjusts by roughly the same amount whether there is conflict or not, explaining its lower QA performance. Additionally, for summarization tasks, AdaCAD generates more faithful summaries whereas other methods tend to hallucinate details.
@@
2 Related Work
Knowledge Conflict.
Integrating external context knowledge into LLMs enables them to keep abreast of current world knowledge (Kasai et al., 2023), reduce hallucination (Shuster et al., 2021), and generate more factual and accurate response. However, a recent line of work focuses on discrepancies between external contextual knowledge and the model’s parametric knowledge, such as LLMs’ over-reliance on their parametric knowledge on entity-based QA tasks Longpre et al. (2021), ignoring retrieved contexts Tan et al. (2024), and exhibiting confirmation bias Xie et al. (2024), etc. More broadly, knowledge conflict relates to the problem of belief revision in LLMs; see Hase et al. (2024) for a more thorough discussion. Other work has proposed solutions for mitigating conflict. Zhou et al. (2023) demonstrate that LLMs’ faithfulness to the context can be significantly improved using carefully designed prompting strategies – this is orthogonal to our work, which is compatible with different prompts. Zhang et al. (2023) address how to combine retrieved and parametric knowledge to improve open-domain QA, but require further training discriminators with silver labels, whereas our method is training-free.
Contrast in Text Generation.
Contrastive approaches for text generation have been widely studied and used to enhance response diversity in conversations Li et al. (2016), steering model generations towards desired attributes while maintaining fluency and diversity Liu et al. (2021), and contrasting between larger and relatively smaller language models to generate high-quality text Li et al. (2023), and improve visually-grounded generation tasks (Wan et al., 2024). Context-aware decoding (CAD; Shi et al., 2024) leverages a contrastive output distribution that amplifies the differences between the output probabilities predicted by a model with and without the context, promoting greater attention to the input context for more faithful and reliable text generation. Unlike AdaCAD, these past contrastive approaches do not adapt the weight on distributions to varying degrees of knowledge conflict. To address this, Yuan et al. (2024) introduce COIECD, a decoding-time method that categorizes instances into two discrete bins – high and low conflict – based on a complex information-entropy constraint governed by tuned hyperparameters, and employs different decoding strategies (by altering CAD) for each. In contrast, AdaCAD employs a single dynamic instance-level strategy that automatically models (based on Jensen-Shannon divergence) a continuous degree of conflict without imposing rigid categories. In addition to these conceptual differences, in Section 4, we show that AdaCAD outperforms both CAD and COIECD on QA and summarization.
@@ 3.1 Task and Notation
Given an input query $\boldsymbol{x}$ with a relevant context $\boldsymbol{c}$, a language model parameterized by $\theta$ is tasked with generating a correct response $\boldsymbol{y}=y_{1},\ldots,y_{n}$ of length $n$ that respects the context. At each decoding step $t$, a token $y_{t}$ can be sampled autoregressively from a probability distribution conditioned on query $\boldsymbol{x}$ and context $\boldsymbol{c}$ as $y_{t}\sim p_{\theta}(y_{t}\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})$. However, when there is conflict between knowledge in the context $\boldsymbol{c}$ and parametric knowledge encoded in LLM, the model can struggle to pay enough attention to $\boldsymbol{c}$ and overly rely on the parametric knowledge (Longpre et al., 2021; Chen et al., 2022), i.e., sample from a distribution more akin to $p_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{<t})$.
3.2 Background: Context-aware Decoding
@@
To mitigate knowledge conflicts, Shi et al. (2024) introduce Context-aware Decoding (CAD), which samples from a contrastive output distribution that amplifies the difference between output probabilities with and without context. CAD measures the parametric knowledge via $p_{\theta}(y_{t}|\boldsymbol{x},\boldsymbol{y}_{t})$ and prioritizes relevant contextual knowledge over the model’s parametric knowledge by using the pointwise mutual information (PMI) between the context $\boldsymbol{c}$ and the generation $y_{t}$, conditioned on $\boldsymbol{x},\boldsymbol{y}_{<t}$ to modify the model’s original output distribution.
$y_{t}$ $\sim\tilde{p}_{\theta}\left(y_{t}|\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t}\right)$
$\propto p_{\theta}(y_{t}\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})\!\left[\frac{p_{\theta}(y_{t}\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})}{p_{\theta}(y_{t}\mid\boldsymbol{x},\boldsymbol{y}_{<t})}\right]^{\alpha}$ (1)
@@
where the PMI term $\frac{p_{\theta}\left(y_{t}\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t}\right)}{p_{\theta}\left(y_{t}\mid\boldsymbol{x},\boldsymbol{y}_{<t}\right)}$ is a scaling factor used to adjust the parametric knowledge, and $\alpha$ is a hyperparameter governing the weight or degree of adjustment. A larger $\alpha$ means a greater adjustment and $\alpha=0$ reduces to no adjustment, i.e., greedy decoding. [^2] While PMI also measures the amount of conflict between the distributions with and without context, empirically, we find that it still results in a high degree of perturbation to the output distribution in cases of low conflict (c.f. Section 5.2).
3.3 AdaCAD: Handling Variable Conflict
@@
In test-time contrastive methods (e.g., Li et al., 2023; Shi et al., 2024), $\alpha$ is a fixed hyperparameter set for an entire dataset, requiring tuning on a validation set. However, every instance in the dataset may need a different weight for adjustment; furthermore, in longer-form generation, individual timesteps may in fact require different weights, making a single $\alpha$ value suboptimal. For instance, in the presence of a high degree of conflict, e.g., Fig. 2 (top), a larger $\alpha$ can perturb the LLM’s output distribution to mitigate over-reliance on parametric knowledge, whereas in cases with low or no conflict (as in Fig. 2 (bottom)), the adjustment to the LLM’s output distribution is minimal. Therefore, a fixed $\alpha$ may fail on scenarios where there are heterogeneous examples with and without conflict, i.e., on realistic datasets.
@@
To address variable conflict, we introduce a different $\alpha_{t}$ for each timestep and each instance. Specifically, we automatically infer $\alpha_{t}$ dynamically based on the degree of knowledge conflict for each instance (and decoding step) without supervision, enabling automatic adaptation. To accomplish this, we use Jensen-Shannon divergence (JSD; Lin, 1991) to model the degree of conflict between the context and parametric knowledge. While similar to Kullback-Leibler divergence, JSD is symmetric and bounded within the range $[0,1]$, making it more suitable for modeling conflicts, as it provides a more interpretable and normalized measure of divergence (details in Appendix A). A larger JSD between $p_{\theta}(y_{t}\mid\boldsymbol{x},\boldsymbol{y}_{t})$ and $p_{\theta}(y_{t}\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{t})$ reflects a greater conflict between context and parameter knowledge, suggesting that we need a larger $\alpha$ to encourage the LM to rely more on the context, while a smaller JSD reflects a smaller conflict, suggesting that a smaller $\alpha$ is required to maintain the LM’s adherence to its parametric knowledge. Therefore, we dynamically set the $\alpha^{\textsc{jsd}}_{t}$ at each decoding time step $t$ to:
@@
$y_{t}\sim p_{\theta}(y_{t}\!\mid\!\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})\!\left[\frac{p_{\theta}(y_{t}\!\mid\!\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})}{p_{\theta}(y_{t}\!\mid\!\boldsymbol{x},\boldsymbol{y}_{<t})}\right]^{\!\alpha^{\textsc{jsd}}_{t}}$
@@
3.4 AdaCAD for Long-form Generation
@@ 4.1 Experimental Setup
Datasets and Metrics.
We evaluate on several QA datasets: Natural Question (NQ; Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2023), and HotpotQA (Yang et al., 2018). We use these datasets to simulate real scenarios with varying degrees of conflict for each instance. Additionally, we evaluate on an existing knowledge conflict dataset, NQ-Swap (Longpre et al., 2021), which is based on the NQ dataset and consists of synthetic conflicting data. We also test on tabular question-answering, using the popular TabMWP dataset (Lu et al., 2023) as a representative example. TabMWP requires mathematical reasoning with tabular contexts. We use exact match accuracy to evaluate all QA datasets.
In order to test AdaCAD on longer-form generation tasks, we evaluate performance on three standard summarization datasets. First, we test on two widely-used datasets: CNN-DM (See et al., 2017) and XSum (Narayan et al., 2018). These datasets are older and many of their documents are contained in the pertaining data. [^4] We use pile.dataportraits.org (Marone and Van Durme, 2023) to test whether examples appear in the pretraining data. We find that many documents from CNN-DM appear in the Pile (Gao et al., 2020), an open-source dataset of English text that is widely used to train various LLMs, suggesting the datasets may have been leaked during pretraining. Therefore, testing solely on these datasets may not provide a comprehensive comparison. To address this, we include TofuEval (Tang et al., 2024), a more recent and challenging benchmark on topic-focused dialogue summarization, where past work has found that LLMs make a substantial number of factual errors, especially when prompted to focus on a marginal topic – secondary information in a document that is not the main focus. Following previous work Shi et al. (2024), we use two reference-based metrics, ROUGE-L (Lin, 2004) and BERT-P (Zhang et al., 2020), to evaluate summarization quality. Following Tang et al. (2024), we also use AlignScore (Zha et al., 2023) to measure the factual consistency of summaries. This reference-free metric measures the factual consistency between the input and output of LLMs. It is standard for TofuEval, which does not support reference-based evaluation (Tang et al., 2024). On TofuEval, we report performance on both the main topics, which are central to the document, and marginal topics, which are less thoroughly explored. For additional details and examples from all datasets, refer to Appendix B.
Source of Context.
We use the gold context provided by NQ, NQ-Swap, TriviaQA, and HotpotQA as the relevant contexts. Since PopQA does not provide gold contexts, we employ BM25 (Robertson and Zaragoza, 2009), a popular off-the-shelf retrieval system, to retrieve relevant contexts from Wikipedia. For TabMWP, we take the semi-structured table as the relevant context. Each table contains multiple rows and columns, and each cell can be a string of text, a string of a number, or a mix of them. In summarization tasks, the source document serves as the relevant context, while the instruction is used as the input query. Examples of the input query $\boldsymbol{x}$ and relevant context $\boldsymbol{c}$ for different datasets are shown in Table 9 of Appendix B. The prompts used are provided in Appendix F.
Models.
@@
Table 1: Results on QA datasets under a zero-shot setting with different base language models show that on average (across tasks and models) AdaCAD improves accuracy by $14.21\%$ over CAD and $4.82\%$ over COIECD.
Baselines.
We compare AdaCAD to standard decoding, context-aware decoding (CAD; Shi et al., 2024) – which has a fixed $\alpha$, and COIECD (Yuan et al., 2024) – which classifies whether there is knowledge conflict using a method controlled by tuned thresholds and then operates in two different decoding modes, each with the same fixed $\alpha$. For consistency, we use the same sampling strategies across all methods: greedy decoding for both QA and summarization tasks. [^5] We found that greedy decoding outperforms top-$p$ sampling on CNN-DM, so we use greedy decoding across all methods for summarization tasks instead of top-$p$ sampling as in Shi et al. (2024). The experimental results are shown in Table 8 of Appendix D. For our baselines, we adopt the best settings reported in prior work: for CAD, we set $\alpha=1$ for the QA datasets and $\alpha=0.5$ for the summarization datasets, following prior work Shi et al. (2024). For COIECD, the values of $\lambda$ and $\alpha$ are set to 0.25 and 1 for QA datasets, and 0.25 and 0.5 for the summarization datasets, respectively, following Yuan et al. (2024). In AdaCAD, the $\alpha$ values are dynamically adjusted based on the degree of knowledge conflict for each instance. All generations including baselines and AdaCAD are under a zero-shot setting.
4.2 Main Results
QA Tasks.
We demonstrate experimental results on QA datasets in Table 1, where AdaCAD consistently outperforms greedy decoding, CAD, and COIECD. For instance, on Llama3-70B, AdaCAD achieves an average score improvement of $2.18\%$ (absolute) over greedy decoding, $12.91\%$ over CAD, and $3.52\%$ over COIECD. Notably, while CAD outperforms greedy decoding by a wide margin on NQ-Swap (which contains only high-conflict examples by design), it suffers major performance losses on other QA datasets, leading to an $18.58\%$ accuracy drop in performance on average across all models; in contrast, AdaCAD performs well across datasets, whether they have conflict or not. Furthermore, AdaCAD consistently outperforms COIECD across various QA datasets, highlighting the strength of our continuous JSD-based approach over COIECD’s binary classification approach that splits instances into ones with conflict or without. For instance, AdaCAD outperforms COIECD by a large average margin of $10.29\%$ on NQ-SWAP across all models. Additionally, on more complex datasets like TabMWP with newer LLMs, AdaCAD also shows superior performance against all baselines, e.g., achieving average improvements of $7.20\%$ with Llama3-70B and $9.50\%$ with Mistral-7B. These results indicate that AdaCAD is better able to combine the advantages of greedy decoding and CAD, performing well in scenarios without knowledge conflict (as greedy decoding does) as well as those with conflict (as CAD does). This combination allows AdaCAD to handle naturalistic datasets with heterogeneous knowledge conflict. Additionally, we also test AdaCAD on all QA datasets using instruction-tuned language models, and as shown in Table 6 of Appendix C, AdaCAD achieves comparable or better performance than all baselines.
@@ 4.2 Main Results
Summarization Tasks.

### v2

@@ Abstract
Knowledge conflict arises from discrepancies between information in the context of a large language model (LLM) and the knowledge stored in its parameters. This can hurt performance when using standard decoding techniques, which tend to ignore the context. Existing test-time contrastive methods seek to address this by comparing the LLM’s output distribution with and without the context and adjust the model according to the contrast between them. However, we find that these methods frequently misjudge the degree of conflict and struggle to handle instances that vary in their amount of conflict, with static methods over-adjusting when conflict is absent. We propose a fine-grained, instance-level approach called AdaCAD, which dynamically infers the weight of adjustment based on the degree of conflict, as measured by the Jensen-Shannon divergence between distributions representing contextual and parametric knowledge. Across four LLMs, six question-answering (QA) and three summarization datasets, we demonstrate that AdaCAD consistently outperforms other decoding baselines with average QA accuracy gains of $14.21\%$ (absolute) over a static contrastive baseline, and improves the factuality of summaries by $6.19$ (AlignScore). Lastly, we show that while contrastive baselines hurt performance when conflict is absent, AdaCAD mitigates these losses, making it more applicable to real-world datasets in which some examples have conflict and others do not. [^1] Our code is publicly available at: https://github.com/HanNight/AdaCAD
@@
Figure 1: In cases of high knowledge conflict, greedy decoding fails to attend to the context, resulting in incorrect answers. Contrastive decoding allows the context to be incorporated, but must be done with care: in low-conflict cases, excessive contrast can over-correct (e.g., by CAD with $\alpha=1$), resulting in incorrect outputs. AdaCAD dynamically adjusts the degree of contrast, allowing it to handle both high and low-conflict cases.
@@ 1 Introduction
We demonstrate AdaCAD’s effectiveness on a diverse range of tasks, covering question-answering (QA) and summarization, with six QA datasets (Natural Question (NQ; Kwiatkowski et al., 2019), NQ-Swap (Longpre et al., 2021), TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2023), HotpotQA (Yang et al., 2018), TabMWP (Lu et al., 2023)) and three summarization datasets (CNN-DM (See et al., 2017), XSum (Narayan et al., 2018), and TofuEval (Tang et al., 2024)). We test a range of base LLMs, examining Llama-2 (Touvron et al., 2023), Llama-3 (AI@Meta, 2024), and Mistral (Jiang et al., 2023). We consider different sizes of these models and also test both the base and instruction-tuned variants. Our results and analyses show that decoding with a uniform level of contrast benefits high-conflict scenarios but generally hurts performance, while the adaptive contrast of AdaCAD results in improvements across the board. Overall, AdaCAD generally achieves superior performance compared to the baselines, with an absolute gain of $14.21\%$ over CAD (a static baseline), $4.82\%$ over COIECD (Yuan et al., 2024, a baseline that classifies instances as conflicting or not), $5.86\%$ over ConfCD (Zhao et al., 2024, a method that makes dynamic token-level adjustments based on LLM confidence), and $2.41\%$ over greedy decoding when averaged across models and QA datasets. On summarization, AdaCAD improves summary quality and factuality, with an average AlignScore (Zha et al., 2023) gain of 4.16 over greedy decoding, 2.19 over CAD, 10.44 over COIECD, and 7.96 over ConfCD.
@@
Furthermore, in our analyses, we explore why AdaCAD improves over the baselines. We first validate the hypothesis that AdaCAD is able to balance contextual and parametric knowledge by assigning lower weights to lower-conflict instances, testing each method on datasets designed to have high and low conflict, finding that AdaCAD’s inferred weight is much lower when there is no conflict. We also compare the amount by which CAD and AdaCAD adjust the base model’s distribution on examples with and without conflict, finding that while AdaCAD changes the distribution less when there is no conflict (i.e., when the base model’s distribution is already sufficient), CAD adjusts by roughly the same amount whether there is conflict or not, explaining its lower QA performance. Additionally, for summarization tasks, AdaCAD generates more faithful summaries whereas other methods tend to hallucinate details.
2 Related Work
Knowledge Conflict
Integrating external knowledge as context into LLMs enables them to keep abreast of current world knowledge (Kasai et al., 2023), reduce hallucination (Shuster et al., 2021), and improve factuality. However, a recent line of work focuses on discrepancies between external contextual knowledge and the model’s parametric knowledge, such as LLMs’ over-reliance on their parametric knowledge on entity-based QA tasks Longpre et al. (2021), ignoring retrieved contexts Tan et al. (2024), and exhibiting confirmation bias Xie et al. (2024), etc. Zhou et al. (2023) demonstrate that LLMs’ faithfulness to the context can be significantly improved using carefully designed prompting strategies – this is orthogonal to our work, which is compatible with different prompts. Zhang et al. (2023) address how to combine retrieved and parametric knowledge to improve open-domain QA, but require further training discriminators with silver labels, whereas our method is training-free.
Contrast in Text Generation
Contrastive approaches for text generation have been widely studied and used to enhance response diversity in conversations Li et al. (2016), steering model generations towards desired attributes while maintaining fluency and diversity Liu et al. (2021), and contrasting between larger and relatively smaller language models to generate high-quality text Li et al. (2023), and improve visually-grounded generation tasks (Wan et al., 2024). Context-aware decoding (CAD; Shi et al., 2024) leverages a contrastive output distribution that amplifies the differences between the output probabilities predicted by a model with and without the context, promoting greater attention to the input context for more faithful and reliable text generation. Unlike AdaCAD, these past contrastive approaches do not adapt the weight on distributions to varying degrees of knowledge conflict. To address this, Yuan et al. (2024) introduce COIECD, a decoding-time method that categorizes instances into two discrete bins – high and low conflict – based on a complex information-entropy constraint governed by tuned hyperparameters, and employs different decoding strategies (by altering CAD) for each. Zhao et al. (2024) uses LLM confidence to adjust the output probabilities dynamically (denoted as ConfCD) as well as relies on additional noisy and irrelevant contexts. In contrast, AdaCAD employs a single dynamic instance-level strategy that automatically models (based on Jensen-Shannon divergence) a continuous degree of conflict without imposing rigid categories or requiring additional noisy and irrelevant contexts, accommodating more general knowledge conflict settings. In addition to these conceptual differences, in Section 4.2, we show that AdaCAD outperforms CAD, COIECD, and ConfCD on QA and summarization.
@@ 3 Methodology
Given an input query $\boldsymbol{x}$ with a relevant context $\boldsymbol{c}$, a language model parameterized by $\theta$ is tasked with generating a correct response $\boldsymbol{y}=y_{1},\ldots,y_{n}$ of length $n$ that respects the context. At each decoding step $t$, a token $y_{t}$ can be sampled autoregressively from a probability distribution conditioned on query $\boldsymbol{x}$ and context $\boldsymbol{c}$ as $y\sim p_{\theta}(y\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})$. However, when there is conflict between knowledge in the context $\boldsymbol{c}$ and parametric knowledge encoded in LLM, the model can struggle to pay enough attention to $\boldsymbol{c}$ and overly rely on the parametric knowledge (Longpre et al., 2021; Chen et al., 2022), i.e., sample from a distribution more akin to $p_{\theta}(y|\boldsymbol{x},\boldsymbol{y}_{<t})$.
Background: Context-aware Decoding
@@
To mitigate knowledge conflicts, Shi et al. (2024) introduce Context-aware Decoding (CAD), which samples from a contrastive output distribution that amplifies the difference between output probabilities with and without context. CAD measures the parametric knowledge via $p_{\theta}(y|\boldsymbol{x},\boldsymbol{y}_{t})$ and prioritizes relevant contextual knowledge over the model’s parametric knowledge by using the pointwise mutual information (PMI) between the context $\boldsymbol{c}$ and the generation $y$, conditioned on $\boldsymbol{x},\boldsymbol{y}_{<t}$ to modify the model’s original output distribution.
$y_{t}$ $\sim\tilde{p}_{\theta}\left(y\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t}\right)$
$\propto p_{\theta}(y\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})\!\left[\frac{p_{\theta}(y\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})}{p_{\theta}(y\mid\boldsymbol{x},\boldsymbol{y}_{<t})}\right]^{\alpha}$ (1)
@@
where the PMI term $\frac{p_{\theta}\left(y\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t}\right)}{p_{\theta}\left(y\mid\boldsymbol{x},\boldsymbol{y}_{<t}\right)}$ is a scaling factor used to adjust the parametric knowledge, and $\alpha$ governs the weight or degree of adjustment. A larger $\alpha$ means a greater adjustment and $\alpha=0$ reduces to no adjustment, i.e., greedy decoding. [^2] While PMI also measures the amount of conflict between the distributions with and without context, empirically, we find that it still results in a high degree of perturbation to the output distribution in cases of low conflict (c.f. Section 5.2).
AdaCAD: Handling Variable Conflict
@@
In test-time contrastive methods – such as those presented by Li et al. (2023) and Shi et al. (2024)$\alpha$ is a fixed hyperparameter set for an entire dataset, requiring tuning on a validation set. However, every instance in the dataset may need a different weight for adjustment; furthermore, in longer-form generation, individual timesteps may require different weights, making a single $\alpha$ value suboptimal. For instance, in the presence of a high degree of conflict, e.g., Fig. 2 (top), a larger $\alpha$ can perturb the LLM’s output distribution to mitigate over-reliance on parametric knowledge, whereas in cases with low or no conflict (as in Fig. 2 (bottom)), the adjustment to the LLM’s output distribution is minimal. Therefore, a fixed $\alpha$ may fail on scenarios where there are heterogeneous examples with and without conflict, i.e., on realistic datasets.
@@
To address variable conflict, we introduce a different $\alpha_{t}$ for each timestep and each instance. Specifically, we automatically infer $\alpha_{t}$ dynamically based on the degree of knowledge conflict for each instance (and decoding step) without supervision, enabling automatic adaptation. To accomplish this, we use Jensen-Shannon divergence (JSD; Lin, 1991) to model the degree of conflict between the context and parametric knowledge. While similar to Kullback-Leibler divergence, JSD is symmetric and bounded within the range $[0,1]$, making it more suitable for modeling conflicts, as it provides a more interpretable and normalized measure of divergence (details in Appendix A). A larger JSD between $p_{\theta}(y\mid\boldsymbol{x},\boldsymbol{y}_{t})$ and $p_{\theta}(y\mid\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{t})$ reflects a greater conflict between context and parameter knowledge, suggesting that we need a larger $\alpha$ to encourage the LM to rely more on the context, while a smaller JSD reflects a smaller conflict, suggesting that a smaller $\alpha$ is required to maintain the LM’s adherence to its parametric knowledge. Therefore, we set $\alpha^{\textsc{jsd}}_{t}$ at each decoding step $t$ to:
@@
$y_{t}\sim p_{\theta}(y\!\mid\!\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})\!\left[\frac{p_{\theta}(y\!\mid\!\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})}{p_{\theta}(y\!\mid\!\boldsymbol{x},\boldsymbol{y}_{<t})}\right]^{\!\alpha^{\textsc{jsd}}_{t}}$
@@
AdaCAD for Long-form Generation
@@ 4.1 Experimental Setup
Datasets and Metrics
We evaluate on several QA datasets: Natural Questions (NQ; Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), PopQA (Mallen et al., 2023), and HotpotQA (Yang et al., 2018). We use these datasets to simulate real scenarios with varying degrees of conflict for each instance. Additionally, we evaluate on an existing knowledge conflict dataset, NQ-Swap (Longpre et al., 2021), which is based on the NQ dataset and consists of synthetic conflicting data. Lastly, we also test on a popular tabular question-answering dataset, TabMWP (Lu et al., 2023), that requires LLMs to use reasoning skills over tabular contexts. We report exact match accuracy on all QA datasets.
@@
Table 1: Under a zero-shot setting, we show that on average (across tasks and models) AdaCAD improves accuracy by $14.21\%$ over CAD, $4.82\%$ over COIECD, and $5.86\%$ over ConfCD (results with instruction-tuned models in Appendix C).
To test AdaCAD on longer-form generation tasks, we evaluate on three standard summarization tasks: CNN-DM (See et al., 2017), XSum (Narayan et al., 2018), and TofuEval (Tang et al., 2024). While many documents from older datasets (such as CNN-DM and XSum) are present in LLM’s pretraining data,444 Using pile.dataportraits.org, we find several documents from CNN-DM appear in the Pile (Gao et al., 2020), commonly used to pretrain LLMs. TofuEval is a recent, more challenging benchmark on topic-focused dialogue summarization (especially for marginal or secondary topics in the document). We use two reference-based metrics, ROUGE-L (Lin, 2004) and BERT-P (Zhang et al., 2020), to evaluate summarization quality. As TofuEval does not support reference-based evaluation Tang et al. (2024), we use recommended AlignScore (Zha et al., 2023) to measure the factual consistency of summaries on both main (central to the document) and marginal (lesser explored) topics. For additional details and examples of all datasets, refer to Appendix B.
Source of Context
We use the gold context provided by NQ, NQ-Swap, TriviaQA, and HotpotQA as the relevant contexts. Since PopQA does not provide gold contexts, we employ BM25 (Robertson and Zaragoza, 2009), to retrieve relevant contexts from Wikipedia. For TabMWP, we take the semi-structured table as the relevant context. In summarization tasks, the source document serves as the relevant context, while the instruction is used as the input query. A summary of input query $\boldsymbol{x}$ and context $\boldsymbol{c}$ for all datasets is shown in Table 10 with corresponding prompts in Appendix F.
Models
@@
Baselines
We compare AdaCAD to standard decoding, context-aware decoding (CAD; Shi et al., 2024) – which has a fixed $\alpha$, COIECD (Yuan et al., 2024) – which classifies whether there is knowledge conflict using a method controlled by tuned thresholds and then operates in two different decoding modes, each with the same fixed $\alpha$, and ConfCD (Zhao et al., 2024) – which dynamically sets alpha based on LLM confidence. Across all tasks and baselines, we use greedy decoding under a zero-shot setting. [^5] We find that greedy decoding outperforms top-$p$ sampling on CNN-DM, so we use greedy decoding across all methods for summarization tasks instead of top-$p$ sampling as in Shi et al. (2024) (c.f. Table 9, Appendix D). For CAD, we set $\alpha=1$ for the QA datasets and $\alpha=0.5$ for the summarization datasets, following prior work Shi et al. (2024). For COIECD, the values of $\lambda$ and $\alpha$ are set to 0.25 and 1 for QA datasets, and 0.25 and 0.5 for the summarization datasets, respectively, following Yuan et al. (2024). For ConfCD, the $\alpha$ values are set to the maximum token probability with context ($C_{R}=\max_{y^{\prime}\in V}p_{\theta}(y^{\prime}|\boldsymbol{c},\boldsymbol{x},\boldsymbol{y}_{<t})$) if $C_{R}$ exceeds the maximum token probability without context (i.e. $C_{R}>C=\max_{y^{\prime}\in V}p_{\theta}(y^{\prime}|\boldsymbol{x},\boldsymbol{y}_{<t})$); otherwise, it is given by $1-C$. In AdaCAD, the $\alpha$ values are dynamically adjusted based on the degree of knowledge conflict for each instance.
4.2 Main Results
QA Tasks
From Table 1, we observe that AdaCAD consistently outperforms greedy decoding, CAD, COIECD, and ConfCD. For instance, on Llama3-70B, AdaCAD achieves an average score improvement of $2.18\%$ (absolute) over greedy decoding, $12.91\%$ over CAD, $3.52\%$ over COIECD, and $2.44\%$ over ConfCD. Note that while CAD performs quite well on NQ-Swap (containing only high-conflict examples), it often degrades performance (relative to greedy decoding) on other QA datasets, resulting in an $18.58\%$ accuracy drop on average across all models and tasks; in contrast, AdaCAD performs well across datasets, whether they have conflict or not. Furthermore, AdaCAD consistently outperforms COIECD across various QA datasets, highlighting the strength of our continuous JSD-based approach over COIECD’s binary classification approach that splits instances into ones with conflict or without. For instance, AdaCAD outperforms COIECD by a large average margin of $10.29\%$ on NQ-SWAP across all models. Additionally, on more complex datasets like TabMWP with newer LLMs, AdaCAD also shows superior performance against all baselines, e.g., achieving average improvements of $6.30\%$ with Llama3-70B and $9.23\%$ with Mistral-7B. These results indicate that AdaCAD is better able to combine the advantages of greedy decoding and CAD, performing well in scenarios without knowledge conflict (as greedy decoding does) as well as those with conflict (as CAD does).
Summarization Tasks